OpenStudy (anonymous):

Please help!!!!!!!! If you multiply the four-digit number abcd by 4, the order of digits will be reversed. That is , abcd x 4=dcba. The digits a, b,c and d are all different. Find abcd. That means what numbers are abcd equal to?

OpenStudy (anonymous):

Any help?

ganeshie8 (ganeshie8):

A must be 1 or 2

ganeshie8 (ganeshie8):

and D > A

OpenStudy (anonymous):

Show how you found your answer please

hartnn (hartnn):

maybe we start off like this 4000a+400b+40c+4d=1000d+100c+10b+a

hartnn (hartnn):

3999a+390b=60c+996d

hartnn (hartnn):

4d=a

OpenStudy (anonymous):

How did you get those numbers? I am confused.

hartnn (hartnn):

1234 = 1*1000+2*100+3*10+4

OpenStudy (anonymous):

The answer is 2178 x 4 = 8712 Click this link to find how to solve this: http://www.100marks.in/forum/f18/abcd-x-4-%3D-dcba-aptitude-question-what-b-c-d-332/

OpenStudy (anonymous):

Ok thanks

hartnn (hartnn):

Given condition is A B C D *4 -------- D C B A Outcome is a 4 digit number, i.e no carry from A when we multiply this by 4 So, Possible values of A = 1 or 2 Now A can't be 1, because in final result A is in unit place and 1 is not possible when we multiply any number by 4 So, A =2 Now we have 2 B C D *4 --------- D C B 2 Now you can see that D = 8, 2 B C 8 *4 -------- 8 C B 2 Now you can see taht no carry must be there when we multiply B with 4, so possible values of B is 1 or 2 Let us consider B =1, i.e 2 1 C 8 *4 --------- 8 C 1 2 Now only c = ?, you can use equation (2000+100+10C+8)*4 = 8000+100C+10+2 40 C = 432-12 = 420 C = 7 So we have A =2, B = 1, C = 7, D = 8

ganeshie8 (ganeshie8):

good one :)

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

What does i,e stand for?

hartnn (hartnn):

that is = i.e,

OpenStudy (thewalkingdeadrick):

supercalafragalisticesbealadoshis